(1 point) If A = | -1 0 1+i į 1-2 – 1 0 -1 + 2i] 4 , then –4 + 2i] A-1 =
Answer #1

Given : - - A=/ Hi |-a-i o -tail í 4 o -4 +ật a _ Adj (A) we will find det (A) det (A) -6121€ (-4+2°) - 4(0)) -06 +(-1+27°) (/ + -t = (ن) 6 - 5 = | F3 = (1) -|- = = = a = -442 (أ+ - ( ۵) - (اہ+ہا-) دا / dt+1- را = وه / ها - (1+) + + ) - 3 - ا = ) = 3- Adj (A) = transpose Lauira of Cofactor matrin -gi-2 dil it Loo Iti 1-2-1 -a-i 0 0 = /-41-2 -21-2 121-1 0 of 1-41-2 Hie -si-